OpenStudy (anonymous):

Suzette did twice as many push-ups as Mark did. Bonnie did five less push-ups than Mark did. If they did 71 push-ups together. How many did each do?

jigglypuff314 (jigglypuff314):

so if Suzette = S Mark = M and Bonnie = B then you are given the equations 2*S = M B - 5 = M M + B + S = 71

OpenStudy (anonymous):

ohh

jigglypuff314 (jigglypuff314):

first let's try solving for M if 2*S = M then S = M/2 yes? and B - 5 = M would be B = M +5 then in the equation of M + B + S = 71 you can substitute the values you got so M + (M+5) + (M/2) = 71 then solve for M

OpenStudy (anonymous):

2 x S = m?

jigglypuff314 (jigglypuff314):

... that is one equation yes... try finding M when M + (M+5) + (M/2) = 71

OpenStudy (anonymous):

3m + 5 + 2 = 71?

OpenStudy (anonymous):

so it be 3m + 7 = 71?

OpenStudy (anonymous):



jigglypuff314 (jigglypuff314):

(M/2) is M over 2 which can not be directly added to the other two M values